### v1

# Constant-Size Cryptographic Evidence Structures for Regulated AI Workflows

Abstract
Regulated AI workflows—such as those used in clinical trials, medical decision support, and financial compliance—must satisfy strict auditability and integrity requirements. Existing logging and audit-trail mechanisms typically rely on variable-length records, bulky cryptographic transcripts, or ad-hoc database schemas. These designs suffer from metadata leakage, irregular performance characteristics, and weak alignment with formal security notions and regulatory semantics.
This paper introduces constant-size cryptographic evidence structures, a general abstraction for representing verifiable audit evidence for AI workflows in regulated environments. Each evidence item is a fixed-size tuple of cryptographic fields, designed to (i) provide strong binding to workflow events and configurations, (ii) support constant-size storage and uniform verification cost per event, and (iii) compose cleanly with hash-chain and Merkle-based audit constructions. We formalize a simple model of regulated AI workflows, define syntax and algorithms for evidence structures, and articulate security goals such as audit integrity and non-equivocation.
We present a generic hash-and-sign construction that instantiates this abstraction using a collision-resistant hash function and a standard digital signature scheme. We then show how to integrate the construction with hash-chained logs, Merkle-tree anchoring, and (optionally) ted execution environments, and we analyze the asymptotic complexity of evidence generation and verification. Finally, we implement a prototype library and report microbenchmark results on commodity hardware, demonstrating that the per-event overhead of constant-size evidence is small and predictable. The design is informed by industrial experience with regulated AI systems at Codebat Technologies Inc., while the paper focuses on the abstraction, algorithms, and their security and performance characteristics. This work aims to provide a foundation for standardized audit mechanisms in regulated AI, with implications for clinical trial management, pharmaceutical compliance, and medical AI governance.
1 Introduction
AI systems deployed in regulated domains—including clinical trials, medical imaging, drug safety monitoring, and financial compliance—must provide strong guarantees of traceability, integrity, and accountability. Regulators and institutional review boards (IRBs) require that each critical operation (e.g., data access, model inference, randomization, or protocol change) leave a verifiable audit trail that can be inspected months or years later. At the same time, operators must protect sensitive information such as personal health information (PHI), internal model details, or proprietary algorithms.
In practice, most audit trails are built from one or more of the following components:
  • Variable-length application logs (e.g., syslog-style entries),
  • Ad-hoc database schemas updated by application code,
  • Blockchain or distributed-ledger events with free-form payloads.
These mechanisms are flexible but misaligned with the needs of regulated AI workflows. Variable-length records leak metadata (e.g., larger records may correspond to imaging data or long clinical notes), incur non-uniform storage and verification costs, and complicate high-throughput verification. Ledger-based solutions often incur unpredictable gas or transaction fees and push complexity into the smart-contract layer.
Secure logging and audit-trail systems have been studied extensively [schneier1999secure, crosby2009efficient, bellare1997forward], as have hash-chain and Merkle-tree-based data structures [merkle1980protocols, haber1991time]. More recently, blockchain-based transparency logs [laurie2013certificate, nakamoto2008bitcoin] and ted execution environments (TEEs) [costan2016intel, sabt2015ted] have been used to build tamper-evident records for security-critical operations. However, these approaches typically operate on variable-length or application-specific records and do not treat a fixed-size, hardware-friendly evidence representation as a first-class abstraction.
Our approach.
We propose to view audit evidence for regulated AI workflows as constant-size cryptographic structures generated and verified by well-defined algorithms. Each evidence item is a tuple of a fixed number of fields, where each field is a fixed-length bitstring (e.g., the output of a hash function). Conceptually, each item binds to:
  • A well-defined unit of work (e.g., a model invocation or randomization step),
  • The relevant inputs and outputs (via commitments),
  • The execution environment (e.g., an environment identifier or attestation digest),
  • A link into a larger audit structure (e.g., a hash chain or Merkle root),
  • A cryptographic authenticator that binds all fields together.
The central requirement is that the representation size of each evidence item is independent of the size or type of the underlying event, which enables constant-size storage and uniform verification cost per event. Figure 1 illustrates the high-level architecture of the evidence structure system.
[…]
Figure 1: Overview of the evidence workflow from workflow events to immutable storage and verification.
Contributions.
We make the following contributions:
  1. 1.
    We formalize a simple model of regulated AI workflows and define the syntax of constant-size evidence structures, including generation and verification algorithms.
  2. 2.
    We identify design objectives for such structures, including fixed-size representation, strong cryptographic binding, efficient batch verification, and compatibility with hash-chain and Merkle-based audit frameworks.
  3. 3.
    We present a generic hash-and-sign construction that instantiates the abstraction using a collision-resistant hash function and a digital signature scheme, and we give informal security arguments for audit integrity and non-equivocation.
  4. 4.
    We discuss how the construction composes with hash-chained logs, Merkle-tree anchoring, and optional TEE-based environment binding, and we analyze the asymptotic complexity of evidence generation and verification.
  5. 5.
    We implement a prototype library and report microbenchmark results that illustrate the performance characteristics of the abstraction on commodity hardware.
2 Model and Design Objectives
2.1 Workflow Model
We consider a workflow as a sequence (or directed acyclic graph) of events. Each event $E$ has:
  • a unique identifier $\mathsf{id}(E)$,
  • structured metadata $\mathsf{meta}(E)$ (e.g., actor, time, configuration),
  • a set of input references $\mathsf{in}(E)$,
  • a set of output references $\mathsf{out}(E)$.
We write $M(E)$ for the full metadata associated with $E$, including these components. We do not assume any particular domain (e.g., clinical, financial); the abstraction is intended to be domain-agnostic.
2.2 Adversary and Environment
We assume an adversary who may:
  • Control or compromise some application components that generate events,
  • Observe or tamper with storage and network channels used to store and transmit audit data,
  • Attempt to suppress, insert, or modify audit evidence,
  • Attempt to correlate audit metadata with sensitive events.
We assume the existence of standard cryptographic primitives (hash functions, signature schemes) with their usual security properties, and we treat external anchoring mechanisms (e.g., append-only logs or blockchains) as black-box components that can store short digests.
2.3 Design Objectives
We highlight several objectives that guide our abstraction.
Fixed-size representation.
Each evidence item associated with an event $E$ must have the same size, independent of the structure or length of $M(E)$. The number of fields and their bit-lengths are fixed at system setup time.
Cryptographic binding.
An evidence item should cryptographically bind to the salient aspects of $E$, including its inputs, outputs, configuration, and (optionally) execution environment. Any modification to these aspects should require either finding hash collisions or forging signatures to avoid detection.
Efficient batch verification.
Evidence items should be verifiable independently and in parallel, with verification cost that is uniform across items and dominated by a small number of hash computations and signature verifications.
Compatibility with audit structures.
Evidence items should integrate cleanly with existing audit structures such as hash chains and Merkle trees [merkle1980protocols], so that large collections of items can be anchored succinctly.
Regulator-aligned semantics.
The fields in an evidence item should support straightforward mapping to regulatory concepts such as actors, timestamps, model versions, and policy states, without hard-coding any particular regulation.
3 Evidence Structures: Syntax and Algorithms
3.1 Syntax
Let $\lambda$ be a security parameter. A constant-size evidence structure for workflows is defined by:
  • A finite index set $I=\{0,1,\dots,k-1\}$ for some fixed $k\in\mathbb{N}$,
  • A field length parameter $\lambda$,
  • A type of evidence items:
    $\mathsf{Ev}=\{0,1\}^{\lambda}\times\dots\times\{0,1\}^{\lambda}$
    consisting of $k$ fields $(f_{0},\dots,f_{k-1})$, each a $\lambda$-bit string.
The interpretation of $f_{i}$ is application-specific but must be fixed by the system designer. Typical roles include:
  • context commitments (e.g., to workflow identifiers),
  • commitments to inputs or outputs,
  • environment identifiers or attestations,
  • links to other evidence items or audit structures,
  • signature- or key-related material.
3.2 Algorithms
An evidence structure is equipped with the following algorithms:
  • $\mathsf{Setup}(1^{\lambda})\rightarrow\mathsf{pp}$: generates public parameters $\mathsf{pp}$ (e.g., hash function description, field interpretation).
  • $\mathsf{KeyGen}(1^{\lambda})\rightarrow(sk,pk)$: generates signing and verification keys for the entity that produces evidence.
  • $\mathsf{Generate}(\mathsf{pp},sk,E)\rightarrow(\mathsf{ev},\sigma)$: given an event $E$ and secret key $sk$, produces an evidence item $\mathsf{ev}\in\mathsf{Ev}$ and an authenticator $\sigma$ (e.g., a signature).
  • $\mathsf{Verify}(\mathsf{pp},pk,E,\mathsf{ev},\sigma)\rightarrow\{\mathsf{accept},\mathsf{reject}\}$: checks whether $\mathsf{ev}$ and $\sigma$ are valid for event $E$ under public key $pk$.
  • $\mathsf{Link}(\mathsf{pp},S)\rightarrow\ell$: given a set or sequence $S$ of evidence items, produces a link value $\ell$ (e.g., a hash-chain tip or Merkle root).
4 A Generic Hash-and-Sign Construction
4.1 Components
Let $H:\{0,1\}^{*}\rightarrow\{0,1\}^{\lambda}$ be a collision-resistant hash function, and let $(\mathsf{KeyGen}^{\mathsf{sig}},\mathsf{Sign},\mathsf{Verify}^{\mathsf{sig}})$ be a digital signature scheme that is existentially unforgeable under chosen-message attacks (EUF-CMA). We assume that the system designer provides a family of deterministic encoding functions
$\phi_{i}:\mathcal{E}\rightarrow\{0,1\}^{*}$
for each index $i\in I$, where $\mathcal{E}$ is the set of events.
4.2 Algorithms in Pseudocode
Algorithm 1 $\mathsf{Generate}(\mathsf{pp},sk,E)$
1: Parse $\mathsf{pp}$ to obtain $H$, index set II, encodings $\{\phi_{i}\}_{i\in I}$
2:for each $i\in I$ do
3:  $m_{i}\leftarrow\phi_{i}(E)$ {encode event $E$ for field ii}
4:  $f_{i}\leftarrow H(m_{i})$
5:end for
[…]
8:return $(\mathsf{ev},\sigma)$
Algorithm 2 $\mathsf{Verify}(\mathsf{pp},pk,E,\mathsf{ev},\sigma)$
1: Parse $\mathsf{pp}$ to obtain $H$, index set II, encodings $\{\phi_{i}\}_{i\in I}$
2: Parse $\mathsf{ev}$ as $(f_{0},f_{1},\dots,f_{k-1})$
3:for each $i\in I$ do
4:  $m_{i}\leftarrow\phi_{i}(E)$
5:  $f^{\prime}_{i}\leftarrow H(m_{i})$
6:  if $f^{\prime}_{i}\neq f_{i}$ then
7:   return $\mathsf{reject}$
8:  end if
9:end for
10:if $\mathsf{Verify}^{\mathsf{sig}}_{pk}(\mathsf{ev},\sigma)=0$ then
11:  return $\mathsf{reject}$
12:end if
13:return $\mathsf{accept}$
Algorithm 3 $\mathsf{Link}(\mathsf{pp},S)$ via hash chaining
[…]
2: Let $S=(\mathsf{ev}_{1},\dots,\mathsf{ev}_{n})$ be an ordered sequence
3:for $j=1$ to $n$ do
4:  $\ell_{j}\leftarrow H(\ell_{j-1}\parallel\mathsf{ev}_{j})$
5:end for
6:return $\ell_{n}$
4.3 Informal Security Discussion
Assuming $H$ is collision-resistant and the signature scheme is EUF-CMA secure:
Binding to events.
If an adversary can produce two different events $E\neq E^{\prime}$ and a single evidence item $\mathsf{ev}$ such that both pairs $(E,\mathsf{ev})$ and $(E^{\prime},\mathsf{ev})$ verify under the same key, then either:
  • there exists an index $i$ with $H(\phi_{i}(E))=H(\phi_{i}(E^{\prime}))$ and $\phi_{i}(E)\neq\phi_{i}(E^{\prime})$, giving a collision in $H$, or
  • the adversary can forge a signature on $\mathsf{ev}$ without querying $\mathsf{Sign}$ on that message.
Both events contradict our assumptions.
Tamper detection.
If an adversary modifies any field of a stored evidence item $\mathsf{ev}$, then either the recomputed hashes will not match the encodings of the underlying event, or the original signature will fail to verify. Any successful undetected modification therefore implies either a hash collision or a signature forgery.
Chain integrity and non-equivocation.
When evidence items are linked with Algorithm 3 or with a Merkle tree, an adversary who attempts to remove, re-order, or equivocate about items must either:
  • find collisions in the hash function to produce two distinct sequences with the same chain tip or root, or
  • present two sets of items with different tips/roots and claim they correspond to the same anchored value.
In the first case, this contradicts collision resistance; in the second, the inconsistency is observable to an external auditor that tracks anchored values [haber1991time, laurie2013certificate].
5 Complexity and Algorithmic Integration
5.1 Per-event Cost
For a fixed index set $I$ of size $k$, Algorithm 1 performs:
  • $k$ evaluations of $H$ (one per field),
  • one signing operation.
Algorithm 2 performs:
  • $k$ evaluations of $H$,
  • one signature verification.
Thus, the per-event computational complexity is $\Theta(k)$ hash evaluations plus a constant number of public-key operations, independent of event size or domain.
The representation size of each evidence item is exactly $k\cdot\lambda$ bits, which is constant for a fixed system configuration. In practice, $k$ and $\lambda$ can be chosen to align with hardware-friendly word sizes and cache lines, but the abstraction does not rely on any particular choice.
5.2 Summary of Algorithmic Costs
Table 1: Asymptotic cost of core algorithms for a fixed field count $k$.
Algorithm Hash evaluations Public-key operations
$\mathsf{Generate}$ $k$ 11 sign
$\mathsf{Verify}$ $k$ 11 verify
$\mathsf{Link}$ (chain) $n$ 0
$\mathsf{Link}$ (Merkle) $O(n)$ 0
5.3 Workflow Instrumentation
The evidence structure can be viewed as an instrumentation layer around an existing workflow:
  • For each event $E$, the workflow invokes Algorithm 1 to obtain $(\mathsf{ev},\sigma)$ and stores or exports them to an audit subsystem.
  • An external auditor, given a subset of events and their corresponding evidence items, runs Algorithm 2 to check local consistency.
  • Periodically, a batch of evidence items is passed to Algorithm 3 (or a Merkle-based variant) and anchored externally (e.g., in a transparency log).
Because the evidence items have constant size and uniform cost, the overhead of instrumentation is predictable and can be budgeted at design time.
5.4 Composition with TEEs
If events are generated inside TEEs [costan2016intel, sabt2015ted], the encoding functions $\phi_{i}$ can be defined to include attestation outputs or environment identifiers. For example, one field might be $H(\text{measurement}\parallel\text{config})$, where measurement is the TEE’s measurement of the loaded code. The abstraction provides a fixed-size container into which such identifiers can be hashed.
6 Prototype and Microbenchmarks
To complement the abstract analysis, we implemented a prototype library that instantiates the hash-and-sign construction described above and measured its performance on synthetic workloads. The goal of these experiments is to illustrate the behavior of constant-size evidence structures under representative configurations, rather than to optimize any specific deployment.
6.1 Implementation Overview
The prototype is implemented in and uses a standard 256-bit hash function and an Edwards-curve-based signature scheme. The implementation fixes a small constant $k$ (the number of fields) and a field length $\lambda$, but does not rely on any domain-specific interpretation of fields. Events are represented as structured records with randomly generated metadata; encoding functions $\{\phi_{i}\}$ extract and serialize appropriate components of each event before hashing.
6.2 Experimental Setup
All experiments are conducted on a commodity server with the configuration summarized in Table 2. We generate synthetic workloads consisting of $N\in\{10^{4},10^{5},10^{6}\}$ events. Each event produces exactly one constant-size evidence item and a signature.
Table 2: Experimental hardware configuration.
Component Specification
CPU 16-core server-class processor @ 3.0 GHz
RAM 64 GB
GPU Mid-range discrete GPU (8 GB VRAM)
Storage NVMe SSD
OS 64-bit Linux
Compiler Optimizing native compiler (release build)
6.3 Per-event Generation Cost
Table 3 reports generation throughput and average per-event latency for CPU-only execution, using a representative configuration of $k$ and $\lambda$. We report results for a single-threaded implementation and for a multi-threaded implementation that processes events in parallel across CPU cores.
Table 3: Evidence generation throughput and latency on CPU (synthetic workloads).
[…]
Single-threaded $3.5\times 10^{4}$ 28.4 $k$ fixed, $\lambda$ fixed
Multi-threaded (16 cores) $2.8\times 10^{5}$ 5.7 near-linear scaling
The results are consistent with the complexity analysis: for fixed $k$ and $\lambda$, per-event latency remains in the tens of microseconds, and parallelization across cores yields nearly linear speedup until memory bandwidth becomes a bottleneck.
6.4 Batch Verification on CPU and GPU
We evaluate batch verification performance for $M\in\{10^{4},10^{5},10^{6}\}$ evidence items. Table 4 summarizes the results on CPU and GPU. On the GPU, we launch one thread per evidence item and arrange items in a contiguous buffer to encourage coalesced memory accesses.
Table 4: Batch verification throughput on CPU and GPU (synthetic workloads).
[…]
CPU (16 threads) $2.5\times 10^{5}$ $10^{5}$ 6.1
GPU (mid-range) $4.0\times 10^{5}$ $10^{6}$ 2.5
The fixed-size layout of evidence items leads to regular memory access patterns and uniform control flow, which are favorable for both vectorized CPU execution and GPU kernels.
6.5 Storage Considerations
Because each event produces exactly one constant-size evidence item and a signature, the total storage needed for a workload of $N$ events is linear in $N$, with a constant factor determined by $k$, $\lambda$, and the signature size. For a fixed configuration, storing $10^{6}$ evidence items (plus signatures and minimal metadata) requires on the order of a few gigabytes of disk space. This is comparable to or smaller than the space required for verbose text-based logs for similar workloads, while providing stronger integrity and verification properties.
7 Regulatory Alignment (High-Level)
Although we do not model any particular regulation in detail, constant-size evidence structures naturally support several common requirements:
Audit controls.
Every security-relevant event is associated with a verifiable evidence item, providing a tamper-evident trail [hipaa].
Integrity and origin.
Cryptographic bindings via hashes and signatures provide strong integrity guarantees, and public keys can be mapped to organizational identities to support accountability.
Lifecycle and configuration tracking.
Fields that encode model identifiers, policy states, or configuration hashes support long-term tracking of model versions and configuration changes [muniswamy2006provenance, sultana2014secure].
8 Related Work
Secure logging and audit trails.
Schneier and Kelsey [schneier1999secure] and Bellare and Yee [bellare1997forward] proposed early designs for secure audit logs and forward-secure logging, focusing on cryptographic protection of log entries. Crosby and Wallach [crosby2009efficient] studied efficient data structures for tamper-evident logging. Our evidence structures build on these ideas but introduce a uniform, constant-size representation as a first-class abstraction tailored to regulated AI workflows.
Blockchain and transparency logs.
Merkle trees [merkle1980protocols] and time-stamping systems [haber1991time] underlie modern blockchain and transparency-log designs [nakamoto2008bitcoin, laurie2013certificate]. Our abstraction is designed to be ledger-friendly: evidence items are small and fixed-size, and our $\mathsf{Link}$ algorithm can produce succinct anchors suitable for external logs.
Ted execution environments.
TEEs such as Intel SGX [costan2016intel] and ARM TZone [sabt2015ted] provide hardware-isolated execution for security-critical code. Several systems use TEEs to protect data-processing pipelines and logging [hunt2018ryoan, ohrimenko2016oblivious]. Our design treats TEEs as optional components that can contribute identifiers or attestation digests to the encoding functions $\phi_{i}$, without changing the core evidence abstraction.
Data provenance and regulated AI.
Data provenance systems [muniswamy2006provenance, sultana2014secure] capture the lineage of data transformations, often at the file or database level. In regulated AI settings, provenance must be coupled with cryptographic integrity guarantees and regulator-aligned semantics. Constant-size evidence structures provide a compact, verifiable representation of key events and configurations that can coexist with richer provenance graphs.
Comparison to existing systems.
While secure logging systems [schneier1999secure, crosby2009efficient] and blockchain-based audits [laurie2013certificate, nakamoto2008bitcoin] provide integrity guarantees, they typically optimize for flexibility rather than uniformity. Variable-length records dominate existing frameworks, leading to:
  • Metadata leakage risks in privacy-sensitive domains, where record length may correlate with the type or sensitivity of underlying events,
  • Unpredictable verification costs in high-throughput settings, as parsing and validating records of varying sizes complicates batch processing and introduces irregular memory access patterns,
  • Complex integration with hardware accelerators (GPUs, TEEs), which benefit from uniform data structures and predictable control flow.
Our constant-size abstraction addresses these gaps by treating uniformity as a first-class design constraint, enabling new deployment patterns for regulated AI workflows. To the best of our knowledge, prior work does not expose a constant-size, hardware-friendly evidence abstraction as a primary design goal, nor does it explicitly target the combination of fixed-size evidence, audit composition, and regulator-aligned semantics that we present here.
9 Discussion and Limitations
The proposed abstraction and prototype have several limitations:
  • We concentrate on the evidence structure and its cryptographic properties, rather than on key management, incident response, or organizational processes.
  • Our security discussion is informal and follows established lines from secure logging and transparency systems; a full game-based formalization remains an interesting direction.
  • The microbenchmarks use synthetic workloads and generic encodings; domain-specific deployments may exhibit different performance characteristics.
  • We leave the choice of encoding functions $\phi_{i}$ to system designers; poor choices could undermine privacy or interpretability, even if the cryptographic layer is sound.
Despite these limitations, constant-size cryptographic evidence structures provide a useful building block for regulated AI workflows and can be instantiated in a variety of architectures, including industrial systems.
9.1 Lessons from Industrial Deployment
Our industrial experience deploying variants of this abstraction in regulated environments has revealed several practical considerations that may inform future implementations:
Field allocation.
Determining the optimal number of fields $k$ requires balancing expressiveness against storage cost. In practice, we have found that $k\in[8,12]$ is often sufficient for most clinical and pharmaceutical workflows, as it provides enough capacity to commit to workflow context, inputs, outputs, environment identifiers, model versions, policy states, audit links, and application-specific extensions, without introducing excessive overhead.
Encoding design.
The choice of encoding functions $\phi_{i}$ is critical and domain-dependent. Domain experts (e.g., clinical researchers, compliance officers) must be involved early in the design process to ensure that regulatory concepts (such as informed consent, randomization integrity, or data provenance) map cleanly to evidence fields. A poorly chosen encoding can lead to ambiguity during audits or to inadvertent disclosure of sensitive information.
Anchoring frequency.
[…]
Integration with legacy systems.
Retrofitting constant-size evidence structures into existing workflows can be challenging, particularly when legacy systems produce variable-length logs or lack well-defined event boundaries. In such cases, an intermediate translation layer may be needed to extract structured events and construct evidence items on-the-fly. This translation introduces a t boundary that must be carefully managed.
These insights, while derived from proprietary deployments, are abstracted here to inform future research and to help practitioners anticipate common pitfalls when adopting constant-size evidence abstractions.
Acknowledgments
The author thanks collaborators and colleagues at Codebat Technologies Inc. for discussions on regulated AI workflows and auditability, and for feedback that motivated a focus on abstraction, algorithm design, and performance evaluation. Any errors or omissions are the author’s own.
Certain implementation details are covered by patent applications by Codebat Technologies Inc.; the present paper concentrates on the abstract construction and its security/performance properties.
10 Conclusion
We introduced constant-size cryptographic evidence structures as a general abstraction for representing audit-relevant events in regulated AI workflows. By fixing the size and layout of evidence items and defining simple generation, verification, and linking algorithms, we obtain predictable storage and verification costs, strong cryptographic binding, and clean composition with hash-chained and Merkle-based audit frameworks. A generic hash-and-sign construction shows how standard primitives can instantiate this abstraction, and microbenchmarks from a prototype implementation indicate that the per-event overhead is compatible with high-throughput workloads on commodity hardware.
We believe that constant-size evidence structures can serve as a foundation for industry-wide standards in regulated AI audit trails. The abstraction’s emphasis on uniformity, cryptographic rigor, and hardware-friendliness makes it well-suited to emerging requirements in clinical trials, pharmaceutical manufacturing, medical AI, and financial compliance. We encourage practitioners, regulators, and standards bodies to consider fixed-size representations when designing next-generation compliance frameworks. Open discussion of these abstractions—including their security properties, performance trade-offs, and integration patterns—can accelerate the development of tworthy AI systems across healthcare, finance, and other high-stakes domains.
Future work includes formalizing security notions for evidence structures via game-based definitions, exploring richer encoding strategies that preserve privacy while supporting regulatory semantics (e.g., using zero-knowledge proofs or homomorphic commitments), and empirically studying how such abstractions integrate with full-scale AI systems in regulated environments. We also hope to engage with regulatory bodies and industry consortia to refine the design principles and to explore standardization opportunities.